x = 12t²t³ + t, Determine velocity equation and acceleration equation.

1
Expert's answer
2022-01-25T13:58:01-0500

x = 12t²t³ + t

First simplify the equation

x = 12t(²+³) + t = 12t5+t

to find velocity equation differentiate the simplified equation

d/dx= 60t4 + 1

hence velocity, v= 60t4 + 1


to find acceleration equation differentiate the velocity equation

d/dv= 240t3

hence acceleration, a= 240t3